# Topology of Discrete Quantum Feedback Control

### v1

@@ VI Feedback control with symmetry
In this section, we present examples of feedback control for each symmetry class shown in Table 2. On the basis of the intrinsic modular conjugation symmetry (see Sec. V.2), the ten symmetry classes of quantum channels listed in Table 2 can be categorized into three groups according the presence or absense of $K$ symmetry and the sign of $\eta_{K}$. The first group (AI, ${}_{+}$, and ${}_{-}$) includes $K$ symmetry with $\eta_{K}=+1$. Since the modular conjugation symmetry features $K$ symmetry with $\eta_{K}=+1$, these symmetry classes can be realized with quantum channels without additional symmetries or with additional $C$ or $Q$ symmetry. The second group (A, psH, ${}^{\dagger}$, and ${}^{\dagger}$) does not have $K$ symmetry. These classes can be realized in a symmetry sector of a quantun channel with unitary symmetry. The third group (AII, ${}_{+}$, and ${}_{-}$) includes $K$ symmetry with $\eta_{K}=-1$. Such symmetry classes can be realized by imposing $K$ symmetry with $\eta_{K}=-1$ on the second group.
@@ VI.1 Class AI: feedback without symmetry or with particle-hole symmetry
VI.2 Class ${}_{+}$ and ${}_{-}$: feedback with a two-point projective measurement
The symmetry classes that include $K$ symmetry with $\eta_{K}=+1$ are class AI, ${}_{+}$, and ${}_{-}$, where the last two classes include $C$ and $Q$ symmetries. As mentioned in Sec. V, $C$ and $Q$ symmetries respectively involve transposition and Hermitian conjugation, both of which reverse the order of operators. Here we construct quantum channels with $C$ and $Q$ symmetries by using feedback control with two-point projective measurements. Specifically, we consider a scheme with two projective measurements before and after a unitary operation which is described by a CPTP map
@@ VI.2 Class ${}_{+}$ and ${}_{-}$: feedback with a two-point projective measurement
where $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1}):=|(\hat{U}_{\vec{j}_{1},a_{1}})_{\vec{j}_{2},a_{2};\vec{j}_{1},a_{1}}|^{2}$ is the probability of a measurement outcome $(\vec{j}_{2},a_{2})$ being obtained in the projective measurement performed after feedback control with a measurment outcome $(\vec{j}_{1},a_{1})$. Thus, the conditions (63a) and (63c) for $C$ and $Q$ symmetries follow from symmetry of the transition probability $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1})$.
@@
Thus, the Bloch matrix has the following structure
@@
which can be interpreted as reciprocity in the feedback process. The truncated Bloch matrix (and hence the Bloch matrix) with Eq. (80) belongs to the symmetry class ${}_{+}$ in Table 2. In the present case with Eqs. (79) and (80), the truncated Bloch matrix is Hermitian ($X_{\mathrm{trunc}}^{\dagger}(\vec{k})=X_{\mathrm{trunc}}(\vec{k})$), and therefore its eigenvalues are real. More specifically, the truncated Bloch matrix in a 1D system cannot have a nonzero winding number in the complex plane.
Another symmetry class ${}_{-}$ possesses $C$ symmetry with $\eta_{C}=-1$. An example of such symmetry is given by
@@
The condition (84a) is satisfied if
@@
which represents the reciprocity in which the the two spin components flow in opposite directions. In contrast, since $p(\vec{j}_{2},\sigma|\vec{j}_{1},\sigma^{\prime})\geq 0$, the conditions (84b) and (84c) are satisfied if and only if
@@
VI.3 Class A, ${}^{\dagger}$, ${}^{\dagger}$, and psH: feedback with unitary symmetry
@@ VI.3.1 Class A
Let $\{\hat{M}_{m}\}$ be measurement operators. We assume that the measurement operators commute with an observable $\hat{C}$: $[\hat{M}_{m},\hat{C}]=0\ (\forall m)$. Such measurement operators describe a quantum nondemolition (QND) measurement for the observable $\hat{C}$ in the sense that all moments of $\hat{C}$ before and after the measurement stay the same [97], which can be shown as
@@
By using strong symmetry, the CPTP map for the feedback control can be block-diagonalized according to the symmetry eigenvalues for the ket and bra spaces. Let
@@
The BL symmetry classes are examined for each block-diagonalized matrix ~$\tilde{\mathcal{E}}_{c,c^{\prime}}$. Here, an important point is that each block does not necessarily have the modular conjugation symmetry since
@@
In particular, ~$\tilde{\mathcal{E}}_{c,c^{\prime}}$ with $c\neq c^{\prime}$ does not possess the modular conjugation symmetry, but it is related to ~$\tilde{\mathcal{E}}_{c^{\prime},c}$ by this symmetry. Since neither of ~$\tilde{\mathcal{E}}_{c,c^{\prime}}$ nor ~$\tilde{\mathcal{E}}_{c^{\prime},c}$ with $c\neq c^{\prime}$ is constrained by the modular conjugation symmetry, it belongs to class A if it does not have any other symmetry.
VI.3.2 Class ${}^{\dagger}$
The other classes (psH, ${}^{\dagger}$, and ${}^{\dagger}$) without $K$ symmetry include $C$ or $Q$ symmetry. These symmetries are realized with feedback protocols with two-point measurements simialar to those in Sec. VII.2. Specifically, we consider a spin-$1/2$ particle with two sublattice degrees of freedom, where a quantum state at site $\vec{j}$ with sublattice $a=A,B$ and spin $\sigma=\uparrow,\downarrow$ is denoted by |$\ket{\vec{j},a,\sigma}$. We perform a projective position measurement on this particle, and assume that this measurement does not disturb the spin state of the particle. Thus, a measurement outcome is specified only by the site index $\vec{j}$ and the sublattice index $a$, and the measurement operator is given by
@@ VI.3.2 Class ${}^{\dagger}$
This condition is interpreted as reciprocity in the feedback process. Since the truncated Bloch matrix $X_{\mathrm{trunc}}^{(\uparrow\downarrow)}(\vec{k})$ does not have the modular conjugation symmetry, it belongs to the symmetry class ${}^{\dagger}$.
VI.3.3 Class ${}^{\dagger}$
The symmetry class ${}^{\dagger}$ involves $C$ symmetry with $\eta_{C}=-1$. Provided that
@@ VI.3.4 Class psH
VI.4 Class AII, ${}_{+}$, and ${}_{-}$: feedback with $K$ symmetry and $\eta_{K}=-1$
@@
If we only impose the symmetry (116), the truncated Bloch matrix belongs to class AII. If we further impose $C$ symmetry by using the sublattice degrees of freedom as in Sec. VI.3, we can also realize the truncated Bloch matrix in class ${}_{+}$ and class ${}_{-}$.
@@ VII Topological Maxwell’s demons
In this section, builiding on the symmetry classification of quantum channels developed in Sec. V, we present various models of topological feedback control. Speficially, we focus on quantum feedback control of a particle in a 1D lattice and construct topological quantum channels protected by their point-gap topology and symmetry. According to the general topological classification of non-Hermitian operators in Ref. [34], a non-Hermitian operator in a 1D system can host nontrivial point-gap topology in class A, AI, ${}_{-}$, ${}^{\dagger}$, and AII. In Sec. VII.1, we consider topological feedback control in class AI, where we can combine particle-hole symmetry with the modular conjugation symmmetry to construct a unitary symmetry. In Sec. VII.2, we construct a model of topological feedback control in class ${}_{-}$ and demonstrate that it shows helical spin transport protected by $C$ symmetry. Section VII.3 is devoted to topological feedback control in class A, where strong unitary symmetry enables us to block-diagonalize a CPTP map. In Sec. VII.4, we impose strong unitary symmetry and $C$ symmetry on a CPTP map to construct topological feedback control in class ${}^{\dagger}$, which exhibits the $\mathbb{Z}_{2}$ skin effect.
@@ VII.1 Su-Schrieffer-Heeger demon
First, we present a model of topological feedback control with $K$ symmetry with $\eta_{K}=+1$ that is distinct from the modular conjugation symmetry. This model provides a concrete example for the discussion presented in Sec. VI.1. We consider a 1D lattice with two sublattices. Each unit cell is labeled by an index $j$ and each sublattice is labeled by $a=A,B$. We first perform projective position measurement which completely resolves the sublattice structure. A measurement outcome is specified by $m=(j,a)$, and the corresponding measurement operator is given by $\hat{M}_{m}=\ket{j,a}\bra{j,a}$ |. Then, we perform feedback control using Hamiltonian $\hat{H}_{a}\ (a=A,B)$ that depends only on the measurement outcome of the sublattice index. The feedback Hamiltonians are given by the Su-Schrieffer-Heeger (SSH) model [98]
@@
$-J_{2}\sum_{i}(\ket{j+1,A}\bra{j,B}+\mathrm{H.c.}),$ (121)
$\hat{H}_{B}=$ $-J_{1}^{\prime}\sum_{i}(\ket{j,A}\bra{j,B}+\mathrm{H.c.})$
@@
In each eigenspace of the operator $(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z})\tilde{S}$, modular conjugation symmetry (57) and particle-hole symmetry (127) are equivalent to each other since $(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z})=\pm\tilde{S}$. Thus, the topological characterization of a CPTP map in those eigenspaces reduces to that of a CPTP map with a single $K$ symmetry. Thus, each symmetry sector of the CPTP map belongs to class AI.
@@
Figure 12: Eigenspectrum of a quantum channel for the SSH demon. Blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC in the infinite-size limit, and orange dots constitute the eigenspectrum under the OBC with the number of unit cells $L=30$. The hopping amplitudes are set to $J_{1}=1,J_{2}=0.5,J_{1}^{\prime}=0.4,J_{2}^{\prime}=0.4$. The duration of feedback is $\tau=1$.
@@ VII.2 Helical Maxwell’s demon
Next, we construct a model of symmetry-protected topological feedback control that belongs to class ${}_{-}$. We consider a spin-$1/2$ particle on a 1D lattice with $L$ sites. The Hilbert space of this system is spanned by a basis {|$\{\ket{j,\sigma}\}$ ($j=1,\cdots,L$, $\sigma=\uparrow,\downarrow$). We consider feedback control of this particle with a two-point measurement scheme (see Fig. 13 for a schematic illustration). We first perform a projective measurement of the position and spin with the projection operators $\hat{P}_{j,\sigma}=\ket{j,\sigma}\bra{j,\sigma}$ |. Let $m_{1}=(j_{1},\sigma_{1})$ be an outcome of this measurement. If we obtain a measurement outcome $(j_{1},\uparrow)$, we raise a spin-independent potential at site $j_{1}-1$ to prevent the particle from going to the left. If we obtain a measurement outcome $(j_{1},\downarrow)$, we raise a spin-independent potential at site $j_{1}+1$ to prevent the particle from going to the right. The feedback Hamiltonians are given by
@@
which corresponds to $C$ symmetry (60a) of the CPTP map with $\eta_{C}=-1$, since
@@
hold due to the symmetry of the feedback protocol. Thus, the Bloch matrix belongs to class ${}_{-}$.
From the block-diagonal structure (136), a topological invariant of the quantum channel for this system is given by
$w_{\mathrm{helical}}(\xi_{\mathrm{PG}}):=$ $\frac{1}{2}\Biggl{[}\int_{-\pi}^{\pi}\frac{dk}{2\pi i}\partial_{k}\ln[\xi_{\uparrow\uparrow}(k)+\xi_{\uparrow\downarrow}(k)-\xi_{\mathrm{PG}}]$
@@
where $\xi_{\mathrm{PG}}$ denotes the location of a point gap and we use Eqs. (139a) and (139b) in deriving the second equality. In Fig. 14(a), we show the eigenspectra of the CPTP map under the PBC and the OBC. Each eigenvalue is doubly degenerate due to the Kramers degeneracy ensured by the $C$ symmetry with $\eta_{C}=-1$ [34]. The loop structue of the eigenspectrum under the PBC leads to the winding number $w_{\mathrm{helical}}(\xi_{\mathrm{PG}})=-1$ for a point gap $\xi_{\mathrm{PG}}$ inside the loop. Reflecting the nontrivial point-gap topology, the eigenspectrum under the OBC is drastically changed from that under the PBC, and the right eigenmodes of the spin-up (spin-down) component are localized near the right (left) edges of the system. This is a manifestation of the symmetry-protected non-Hermitian skin effect [72], where the Kramers doublets are localized near the oppsite edges. Physically, the symmetry-protected non-Hermitian skin effect is a consequence of the helical spin transport due to feedback control. Thus, the helical Maxwell’s demon considered here achieves robust feedback-assisted spin transport protected by nontrivial topology of a quantum channel.
@@
Figure 14: (a) Eigenspectrum of a quantum channel for the helical Maxwell’s demon with additional feedback performed after the second measurement. (b) Eigenspectrum of a quantum channel for the helical Maxwell’s demon with no further feedback performed after the second measurement. An enlarged eigenspectrum near the steady state is also shown. Blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC, orange dots constitute the eigenspectrum under the OBC with $L=30$, and green dots constitute that under the OBC with $L=20$. The hopping amplitude is set to $J=1$ and the transverse magnetic field is set to $h_{x}=0.2$. The duration of feedback is $\tau=1$.

### v2

@@ VI Feedback control with symmetry
In this section, we present examples of feedback control for each symmetry class shown in Table 2. On the basis of the intrinsic modular conjugation symmetry (see Sec. V.2), the ten symmetry classes of quantum channels listed in Table 2 can be categorized into three groups according the presence or absense of $K$ symmetry and the sign of $\eta_{K}$. The first group (AI, AI $+$ psH+, and AI $+$ psH-) includes $K$ symmetry with $\eta_{K}=+1$. Since the modular conjugation symmetry features $K$ symmetry with $\eta_{K}=+1$, these symmetry classes can be realized with quantum channels without additional symmetries or with additional $C$ or $Q$ symmetry. The second group (A, psH, AI, and AII) does not have $K$ symmetry. These classes can be realized in a symmetry sector of a quantun channel with unitary symmetry. The third group (AII, AII $+$ psH+, and AII $+$ psH-) includes $K$ symmetry with $\eta_{K}=-1$. Such symmetry classes can be realized by imposing $K$ symmetry with $\eta_{K}=-1$ on the second group.
@@ VI.1 Class AI: feedback without symmetry or with particle-hole symmetry
VI.2 Class AI $+$ psH+ and AI $+$ psH-: feedback with a two-point projective measurement
The symmetry classes that include $K$ symmetry with $\eta_{K}=+1$ are class AI, AI $+$ psH+, and AI $+$ psH-, where the last two classes include $C$ and $Q$ symmetries. As mentioned in Sec. V, $C$ and $Q$ symmetries respectively involve transposition and Hermitian conjugation, both of which reverse the order of operators. Here we construct quantum channels with $C$ and $Q$ symmetries by using feedback control with two-point projective measurements. Specifically, we consider a scheme with two projective measurements intervened by a unitary operation. This process is described by a CPTP map
@@ VI.2 Class AI $+$ psH+ and AI $+$ psH-: feedback with a two-point projective measurement
where $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1}):=|(\hat{U}_{\vec{j}_{1},a_{1}})_{\vec{j}_{2},a_{2};\vec{j}_{1},a_{1}}|^{2}$ is the probability of a measurement outcome $(\vec{j}_{2},a_{2})$ being obtained in the projective measurement performed after feedback control associated with a measurment outcome $(\vec{j}_{1},a_{1})$. Thus, the conditions (68a) and (68c) for $C$ and $Q$ symmetries follow from symmetry of the transition probability $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1})$.
@@
Thus, the Bloch matrix has the following structure:
@@
which can be interpreted as reciprocity in the feedback process. The truncated Bloch matrix (and hence the original Bloch matrix) with Eq. (83) belongs to the symmetry class AI $+$ psH+ in Table 2. In the present case with Eqs. (82) and (83), the truncated Bloch matrix is Hermitian ($X_{\mathrm{trunc}}^{\dagger}(\vec{k})=X_{\mathrm{trunc}}(\vec{k})$), and therefore its eigenvalues are real. More specifically, the truncated Bloch matrix in a 1D system cannot have a nonzero winding number.
Another symmetry class AI $+$ psH- possesses $C$ symmetry with $\eta_{C}=-1$. An example of such symmetry is given by
@@
It follows from Eq. (81) that condition (87a) is satisfied if
@@
which represents the reciprocity in which the the two spin components flow in opposite directions. In contrast, since $p(\vec{j}_{2},\sigma|\vec{j}_{1},\sigma^{\prime})\geq 0$, conditions (87b) and (87c) are satisfied if and only if
@@
VI.3 Class A, AI, AII, and psH: feedback with unitary symmetry
@@ VI.3.1 Class A
Let $\{\hat{M}_{m}\}$ be measurement operators. We assume that the measurement operators commute with an observable $\hat{C}$: $[\hat{M}_{m},\hat{C}]=0\ (\forall m)$. Such measurement operators describe a quantum nondemolition (QND) measurement for the observable $\hat{C}$ in the sense that the expected value of every moment of $\hat{C}$ before and after the measurement stays the same [99], that is,
@@
Due to strong symmetry, the CPTP map for the feedback control can be block-diagonalized according to the symmetry eigenvalues for the ket and bra spaces. Let
@@
The BL symmetry classes are examined for each block-diagonalized matrix ~$\tilde{\mathcal{E}}_{c,c^{\prime}}$. A crucial observation here is that each block does not necessarily have the modular conjugation symmetry since
@@
In particular, ~$\tilde{\mathcal{E}}_{c,c^{\prime}}$ with $c\neq c^{\prime}$ does not possess the modular conjugation symmetry, but it is related to ~$\tilde{\mathcal{E}}_{c^{\prime},c}$ by this symmetry. Thus, it belongs to class A if it does not have any other symmetry.
VI.3.2 Class AI
The other classes (psH, AI, and AII) without $K$ symmetry include $C$ or $Q$ symmetry. These symmetries are realized with feedback protocols with two-point measurements similar to those in Sec. VII.2. Specifically, we consider a spin-$1/2$ particle with two sublattice degrees of freedom, where a quantum state at site $\vec{j}$ with sublattice $a=A,B$ and spin $\sigma=\uparrow,\downarrow$ is denoted by |$\ket{\vec{j},a,\sigma}$. We perform a projective position measurement on this particle, and assume that this measurement does not disturb the spin state of the particle. Thus, a measurement outcome is specified only by the site index $\vec{j}$ and the sublattice index $a$, and the measurement operator is given by
@@ VI.3.2 Class AI
This condition is interpreted as reciprocity in the feedback process. Since the truncated Bloch matrix $X_{\mathrm{trunc}}^{(\uparrow\downarrow)}(\vec{k})$ does not have $K$ symmetry, it belongs to the symmetry class AI.
VI.3.3 Class AII
The symmetry class AII involves $C$ symmetry with $\eta_{C}=-1$. Provided that
@@ VI.3.4 Class psH
VI.4 Class AII, AII $+$ psH+, and AII $+$ psH-: feedback with $K$ symmetry and $\eta_{K}=-1$
@@
If we only impose the symmetry (117), the truncated Bloch matrix belongs to class AII. If we further impose $C$ symmetry by using the sublattice degrees of freedom as in Sec. VI.3, we can also realize the truncated Bloch matrix in class AII $+$ psH+ and class AII $+$ psH-.
@@ VII Topological Maxwell’s demons
In this section, builiding on the symmetry classification of quantum channels developed in Sec. V, we present various models of topological feedback control. In Sec. VII.1-VII.4, we focus on quantum feedback control of a particle in a 1D lattice and construct topological quantum channels protected by their point-gap topology and symmetry. According to the general topological classification of non-Hermitian operators in Ref. [34], a non-Hermitian operator in a 1D system can host nontrivial point-gap topology in classes A, AI, AI $+$ psH-, AII, and AII. In Sec. VII.1, we consider topological feedback control in class AI, where we can combine particle-hole symmetry with the modular conjugation symmmetry to construct a unitary symmetry. In Sec. VII.2, we construct a model of topological feedback control in class AI $+$ psH- and demonstrate that it shows helical spin transport protected by $C$ symmetry. In Sec. VII.3, we consider topological feedback control in class A, where strong unitary symmetry enables us to block-diagonalize a CPTP map. In Sec. VII.4, we impose strong unitary symmetry and $C$ symmetry on a CPTP map to construct topological feedback control in class AII, which exhibits the $\mathbb{Z}_{2}$ skin effect. Finally, in Sec. VII.5, we show an example of quantum feedback control with nontrivial line-gap topology in two dimensions.
@@ VII.1 Su-Schrieffer-Heeger demon
First, we present a model of topological feedback control with $K$ symmetry with $\eta_{K}=+1$ that is distinct from the modular conjugation symmetry. This model provides a concrete example for the discussion presented in Sec. VI.1. We consider a 1D lattice with two sublattices. Each unit cell is labeled by an index $j$ and each sublattice is labeled by $a=A,B$. We first perform projective position measurement which completely resolves the sublattice structure. A measurement outcome is specified by $m=(j,a)$, and the corresponding measurement operator is given by $\hat{M}_{m}=\ket{j,a}\bra{j,a}$ |. Then, we perform feedback control with Hamiltonian $\hat{H}_{a}\ (a=A,B)$ that depends only on the measurement outcome of the sublattice index. The feedback Hamiltonians are given by the Su-Schrieffer-Heeger (SSH) model [100]:
@@
$-J_{2}\sum_{j}(\ket{j+1,A}\bra{j,B}+\mathrm{H.c.}),$ (122)
$\hat{H}_{B}=$ $-J_{1}^{\prime}\sum_{j}(\ket{j,A}\bra{j,B}+\mathrm{H.c.})$
@@
In each eigenspace of the operator $(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z})\tilde{S}$, the modular conjugation symmetry (62) and the particle-hole symmetry (128) are equivalent to each other since $(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z})=\pm\tilde{S}$. Thus, the topological characterization of a CPTP map in those eigenspaces reduces to that of a CPTP map with a single $K$ symmetry. Thus, each symmetry sector of the CPTP map belongs to class AI.
@@
Figure 13: Eigenspectrum of a quantum channel for the SSH demon. Blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC in the infinite-size limit, and orange dots constitute the eigenspectrum under the OBC with the number of unit cells $L=30$. The hopping amplitudes are set to $J_{1}=1,J_{2}=0.5,J_{1}^{\prime}=0.4$, and $J_{2}^{\prime}=0.4$. The duration of feedback is $\tau=1$.
@@ VII.2 Helical Maxwell’s demon
Next, we construct a model of symmetry-protected topological feedback control that belongs to class AI $+$ psH-. We consider a spin-$1/2$ particle on a 1D lattice with $L$ sites. The Hilbert space of this system is spanned by a basis {|$\{\ket{j,\sigma}\}$ ($j=1,\cdots,L$, $\sigma=\uparrow,\downarrow$). We consider feedback control of this particle with a two-point measurement scheme (see Fig. 14 for a schematic illustration). We first perform a projective measurement of the position and spin with the projection operators $\hat{P}_{j,\sigma}=\ket{j,\sigma}\bra{j,\sigma}$ |. Let $m_{1}=(j_{1},\sigma_{1})$ be an outcome of this measurement. If we obtain a measurement outcome $(j_{1},\uparrow)$, we raise a spin-independent potential at site $j_{1}-1$ to prevent the particle from going to the left. If we obtain a measurement outcome $(j_{1},\downarrow)$, we raise a spin-independent potential at site $j_{1}+1$ to prevent the particle from going to the right. The feedback Hamiltonians are given by
@@
which corresponds to $C$ symmetry (65a) of the CPTP map with $\eta_{C}=-1$, since the relations
@@
hold due to the symmetry of the feedback protocol. Thus, the Bloch matrix belongs to class AI $+$ psH-.
From the block-diagonal structure (137), a topological invariant of the quantum channel for this system is given by the spin winding number
$w_{\mathrm{spin}}(\xi_{\mathrm{PG}}):=$ $\frac{1}{2}\Biggl{[}\int_{-\pi}^{\pi}\frac{dk}{2\pi i}\partial_{k}\ln[\xi_{\uparrow\uparrow}(k)+\xi_{\uparrow\downarrow}(k)-\xi_{\mathrm{PG}}]$
@@
where $\xi_{\mathrm{PG}}$ denotes the location of a point gap and we use Eqs. (140a) and (140b) in deriving the second equality. In Fig. 15(a), we show the eigenspectra of the CPTP map under the PBC and the OBC. Each eigenvalue is doubly degenerate due to the Kramers degeneracy ensured by the $C$ symmetry with $\eta_{C}=-1$ [34]. The loop structure of the eigenspectrum under the PBC leads to the winding number $w_{\mathrm{spin}}(\xi_{\mathrm{PG}})=-1$ for a point gap $\xi_{\mathrm{PG}}$ inside the loop. Reflecting the nontrivial point-gap topology, the eigenspectrum under the OBC is drastically changed from that under the PBC, and the right eigenmodes of the spin-up (spin-down) component are localized near the right (left) edge of the system. This is a manifestation of the symmetry-protected non-Hermitian skin effect [72], where the Kramers doublets are localized near the opposite edges. Physically, the symmetry-protected non-Hermitian skin effect is a consequence of the helical spin transport due to feedback control. Thus, the helical Maxwell’s demon considered here achieves robust feedback-assisted spin transport protected by nontrivial topology of a quantum channel.
@@
Figure 15: (a) Eigenspectrum of a quantum channel for the helical Maxwell’s demon with additional feedback performed after the second measurement. (b) Eigenspectrum of a quantum channel for the helical Maxwell’s demon with no further feedback performed after the second measurement. An enlarged eigenspectrum near the steady state is also shown. Blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC, and orange (green) dots form the eigenspectrum under the OBC with $L=30$ ($L=20$). The hopping amplitude is set to $J=1$ and the transverse magnetic field is set to $h_{x}=0.2$. The duration of feedback is $\tau=1$.